# Development of ion-beam sputtered silicon nitride thin films for low-noise mirror coatings of gravitational-wave detectors

### v1

@@ Abstract
In order to test the hypothesis of a correlation between the synthesis conditions of the films and their elemental composition and optical and mechanical properties, we varied the voltage, current intensity and composition of the sputtering ion beam, and we performed a broad campaign of characterizations. While the refractive index was found to monotonically depend on the beam voltage and linearly vary with the N/Si ratio, the optical absorption appeared to be strongly sensitive to other factors, as yet unidentified. However, by systematically varying the deposition parameters, an optimal working point was found. Thus we show that the loss angle and extinction coefficient of our thin films can be as low as $(1.0\pm 0.1)\times 10^{-4}$ rad at $\sim$2.8 kHz and $(6.4\pm 0.2)\times 10^{-6}$ at 1064 nm, respectively, after thermal treatment at 900 C. Such loss angle value is the lowest ever measured on this class of thin films.
We then used our silicon nitride thin films to design and produce a multi-material mirror coating showing a thermal noise amplitude of $(10.3\pm 0.2)\times 10^{-18}$ m Hz-1/2 at 100 Hz, which is 25% lower than in current mirror coatings of the Advanced LIGO and Advanced Virgo interferometers, and an optical absorption as low as $(1.9\pm 0.2)$ parts per million at 1064 nm.
@@ 1 Introduction
The current mirror coatings of LIGO, Virgo and KAGRA are thickness-optimized Bragg reflectors [9] deposited by ion-beam sputtering (IBS) at the Laboratoire des Matériaux Avancés (LMA). They are made of stacked alternate thin layers of tantalum pentoxide (Ta2O5, also known as tantala, high refractive index), either plain or mixed with titanium dioxide (TiO2 or titania, high refractive index), and silicon dioxide (SiO2 or silica, low refractive index) [10, 11]. These coatings are already fully compliant with the optical loss requirements of present gravitational-wave interferometers, but still show a noise amplitude which is expected to be a severe sensitivity limitation [1, 2, 12].
Indeed, according to the fluctuation-dissipation theorem [13], the power spectral density of Brownian noise fluctuations at the coatings’ surface is proportional to the rate of elastic energy dissipation within the thin coatings’ layers. This rate, quantified by the loss angle $\varphi$, must hence be reduced as much as possible to decrease coating thermal noise. At room temperature, the loss angle of tantala and tantala-titania layers is about an order of magnitude higher than that of silica layers [10, 14], making it the dominant source of dissipation. So, since the early 2000s, a considerable research effort is being committed to developing alternative high-index coating materials of lower loss angle [15, 16, 17].
@@ 2.1 Samples
The thin films used in this work were deposited in a custom-developed facility at the LMA. We used several sets of deposition parameters, in order to test the hypothesis of a correlation between the synthesis conditions of the thin-film samples and their composition and optical and mechanical properties. Hence we varied the kinetic energy and current intensity of the sputtering particles, as well as the elemental composition of the ion beam impinging on the target. At the same time, for every deposition run, we kept the beam/target and target/substrate holder angles constant and equal to 45 deg, the working pressure during deposition was of the order of $10^{-4}$ mbar, the substrates were heated up to a temperature between 100 and 150 C while rotating at a speed of 3 rpm. Prior to deposition, the base pressure inside the chamber was lower than $2\times 10^{-7}$ mbar.
@@
Table 1: Relevant deposition parameters of IBS SiNx thin-film samples: ion beam voltage $V$ and current $I$, Ar and N2 gas mass flows fed to the ion source, and average deposition rate $\dot{t}$.
@@
Figure 1: Average deposition rate $\dot{t}$ of IBS SiNx thin-film samples, as a function of ion beam settings: voltage $V$ and current $I$, and ratio N2/Ar of gas mass flows to the source.
@@ 2.2 Annealing
High-temperature thermal treatment, or annealing, is known to generally minimize the optical absorption and loss angle of thin-film coatings [10], hence we studied the effect of annealing on the properties of our specific samples.
@@ 2.3.1 Spectroscopic photometry
Samples were deposited on fused-silica substrates (25.4 mm diameter, 6 mm thickness), then measured for transmittance in the 400–1400 nm wavelength range, at normal incidence, using a Cary 7000 spectroscopic photometer by Agilent.
@@ 2.3.3 Photo-thermal deflection (PTD)
Samples for absorption measurements were deposited on fused-silica substrates (25.4 mm diameter, 6 mm thickness).
@@ 2.4.1 Rutherford back-scattering spectrometry (RBS)
At SAFIR, we used samples grown on glassy carbon substrates ($25\times 25$ mm, 0.5 or 1 mm thickness). Relative atomic concentrations of the samples were deduced by integrating the peaks in the acquired spectra with the SIMNRA software [43, 44]. We also used SIMNRA to simulate a small background signal observed in the low-energy region of all spectra, and subtracted the results of those simulations from the oxygen and nitrogen contributions. We modeled that background signal as 0.1 to 0.3% oxygen ambient contamination of the surface of the glassy carbon substrates.
@@ 2.5 Mechanical properties
We estimated the loss angle $\varphi$ of the samples by measuring the frequency and ring-down time of several mechanical resonances [45] of fused-silica disks (50.0 mm diameter, 0.5 or 1.0 mm thickness), before and after coating deposition, in the 2.5-30 kHz range. In order to avoid systematic damping from suspension and ambient pressure, we used a clamp-free, in-vacuum Gentle Nodal Suspension (GeNS) system [46]. More details about our experimental setup and data analysis can be found in our previous works [10, 47].
We derived the volumetric mass density $\rho$ from mass, area and thickness measurements. We measured the mass with an analytical balance, the area was estimated by measuring the diameter of our coated disks with a caliper, and the thickness was known from photometric and ellipsometric measurements.
@@ 2.6 Grazing-incidence X-ray diffraction (GI-XRD) and X-ray reflectivity (XRR)
For both GI-XRD and XRR measurements, samples were grown on single-side polished silicon substrates (76.2 mm diameter, 0.5 mm thickness).
@@ 2.7 Ar desorption
As a consequence of the deposition process, the presence of argon in the samples was highly probable. In order to assess whether such contaminant was desorbed upon annealing, X-ray photoelectron spectroscopy and thermal-desorption spectroscopy measurements were performed.
@@ 2.7.2 Thermal-desorption spectroscopy (TDS)
TDS measurements were carried out by placing the samples in the quartz tube of an MTI GSL100 vacuum furnace. The tube was connected to a Pfeiffer Vacuum QMS200 mass spectrometer, its base pressure was of the order of 10-7 mbar. The furnace heating rate was set at 3 C/min, and surveys were performed from ambient temperature up to the maximum temperature of 850 C allowed by the setup.
@@ 2.8 Coating Brownian thermal noise (CTN)
Samples for CTN measurements were deposited on fused-silica substrates (25.4 mm diameter, 6 mm thickness).
@@ 3.1 Mitigation of water vapor contamination
In the early phases of this work, ever since the very first characterizations were performed, an evident anomaly was observed in the acquired photometric spectra. We interpreted it as a consequence of a gradient of refractive index $n$ along the direction of the samples’ thickness, with $n$ always decreasing from the surface towards the substrate. Under this hypothesis, such gradient resulted as high as 27%.
@@
Thus we performed a series of measurements of Fourier-transform infrared spectroscopy (FTIR), Raman spectroscopy, XPS (also coupled to sputtering along the direction of sample thickness, achieved via focused ion beam etching), RBS and ERDA. All those measurements, not shown here for conciseness, revealed an abundant presence of oxygen and hydrogen within the samples, whose relative atomic concentrations were as high as 34% and 5%, respectively. Incidentally, we were able to verify that the nitrogen-to-silicon ratio N/Si was equal to or less than 1.0 in all those samples, including those deposited using a pure N2 sputtering beam. This clearly indicated the presence of excess silicon, only partly oxidized.
@@ 3.2.1 Refractive index and extinction
We found that $n$ was a monotonic increasing function of the ion beam voltage $V$, in agreement with results found in the literature [18, 27]. At the same time, we observed that the ion beam elemental composition and current $I$ had a limited effect instead. The voltage $V$ also had a noticeable effect on the $k$ coefficient, whose minimum was obtained at 700 V, and the same held for the ion beam composition. The lowest $k$ values were obtained for a mass flow ratio N2/Ar equal to 2. This outcome is in contrast with results reported in the literature, indicating that the optimal ratio would rather be 4 [53]. Regardless, the largest effect on $k$ was obtained by varying the ion beam current $I$. At $V=700$ V, we were able to decrease $k$ by more than an order of magnitude, from $2.1\times 10^{-4}$ to $1.5\times 10^{-5}$, by increasing $I$ up to 195 mA.
Thus a minimum of extinction was achieved with sample 21025 for $V=700$ V, $I=195$ mA and N2/Ar = 2. Since then, we used those same parameters to deposit multiple nominally identical samples and observed repeatable results. For instance, the average extinction coefficient and refractive index of a subset of fifteen samples were $k=(1.5\pm 0.2)\times 10^{-5}$ and $n=2.04\pm 0.01$, respectively, where the uncertainties are standard deviations. After in-air annealing at 900 C for 10 hours, the extinction decreased to $k=(6.4\pm 0.2)\times 10^{-6}$. Compared to the tantala-titania layers in current mirror coatings of gravitational-wave interferometers, which show $n=2.09\pm 0.01$ and $k\sim 10^{-7}$ [10], the refractive index of sample 21025 is only 2% lower. Its extinction, however, while being amongst the lowest ever measured on this class of thin films [22], is still excessively high.
@@ 3.3 Elemental composition
In order to study more thoroughly the conditions that resulted in minimal optical absorption, we focused our attention on samples deposited on glassy carbon substrates using $I=195$ mA, N2/Ar = 2 and varying $V$. Table 4 presents the relative atomic concentrations of samples 22020, 22022 and 22023, as deduced from acquired RBS spectra. Sample 22020 was deposited using $V=700$ V as in sample 21025, which featured the lowest optical absorption. Samples 22022 and 22023 were deposited using $V=500$ V and $V=900$ V, respectively. By way of example, Fig. 5 shows the measured and simulated RBS spectra of samples 22020 and 22023.
@@
Comparative analysis of RBS results of Table 4 first shows that using $I=195$ mA and N2/Ar = 2 settings slightly increased the N/Si ratio, which reached a value of about 1.1. The beam voltage $V$, on the other hand, only appreciably changed for $V=900$ V, when it equalled 1.0. This seems to contradict the results found in the literature, which indicate that it would be possible to determine a range of $V$ values yielding stoichiometric samples [18, 25, 28, 29, 54]. It should be noted however that these results appear to be in sharp disagreement as to the values to be used. For instance, Bouchier et al. found that stoichiometric samples were obtained using $V\leq 500$ V [28, 29, 54], while Huang et al. and Lee et al. found that stoichiometric samples were instead obtained for 700 V $<V<$ 900 V [25] or even $V=1100$ V [18], respectively. Anyway, it remains that, in our specific case, the best deposition conditions identified so far produced sub-stoichiometric, nitrogen-poor samples.
By combining data of Tables 2 and 4 altogether, we observed that the refractive index $n$ at 1064 nm varies linearly with the Si/N ratios as measured by RBS. This is shown on Figure 6. For the deposition conditions under discussion, we found that
@@ 3.3.1 Contaminants
Fig. 7 shows the measured and simulated ERDA spectra of sample 23017, which are representative of those obtained from our other samples. It can be seen there that, at first glance, the measured relative concentration of hydrogen appeared to be highly inhomogeneous across the sample thickness. However, the two peaks observed at $\sim 453$ and $\sim 630$ keV were likely due to ambient contamination at the substrate and coating surface, respectively. Under this hypothesis, the hydrogen content of sample 23017 should be evaluated in the region in between the peaks, hence amounting to 0.4%. If the peaks were to be included in the analysis, the relative concentration would indeed increase, up to 1.2%. The small signal observed below 400 keV was identified as an artifact of the experimental setup and not considered in our analysis.
Figure 7: Measured and simulated ERDA spectra of IBS SiNx thin-film sample 22020 deposited on glassy carbon.
@@ 3.3.2 Ar desorption
Figure 8: Desorption of Ar from IBS SiNx thin-film sample 22027 upon in-vacuum annealing, compared to the background signal of the experimental setup. The heating rate was 3 C/min.

### v2

@@ Abstract
In order to test the hypothesis of a correlation between the synthesis conditions of the films and their elemental composition and optical and mechanical properties, we varied the voltage, current intensity and composition of the sputtering ion beam, and we performed a broad campaign of characterizations. While the refractive index was found to monotonically depend on the beam voltage and linearly vary with the N/Si ratio, the optical absorption appeared to be strongly sensitive to other factors, as yet unidentified. However, by systematically varying the deposition parameters, an optimal working point was found. Thus we show that the loss angle and extinction coefficient of our thin films can be as low as $(1.0\pm 0.1)\times 10^{-4}$ rad at $\sim$2.8 kHz and $(6.4\pm 0.2)\times 10^{-6}$ at 1064 nm, respectively, after thermal treatment at 900 C. To the best of our knowledge, such loss angle value is the lowest ever measured on this class of thin films.
We then used our silicon nitride thin films to design and produce a multi-material mirror coating showing a thermal noise amplitude of $(10.3\pm 0.2)\times 10^{-18}$ m Hz-1/2 at 100 Hz, which is 25% lower than in current mirror coatings of the Advanced LIGO and Advanced Virgo interferometers, and an optical absorption as low as $(1.6\pm 0.5)$ parts per million at 1064 nm.
@@ 1 Introduction
The current mirror coatings of LIGO, Virgo and KAGRA are thickness-optimized Bragg reflectors [9] deposited by ion-beam sputtering (IBS) at the Laboratoire des Matériaux Avancés (LMA). They are made of stacked alternate thin layers of tantalum pentoxide (Ta2O5, also known as tantala, high refractive index), either plain or mixed with titanium dioxide (TiO2 or titania, high refractive index), and silicon dioxide (SiO2 or silica, low refractive index) [10, 11]. These coatings are already fully compliant with the optical loss requirements of present gravitational-wave interferometers, but still show a thermal noise amplitude which is expected to be a severe sensitivity limitation [1, 2, 12]. For reference, the current Bragg reflectors of LIGO and Virgo feature an optical loss due to absorption and scattering of 1 and 5 parts per million (ppm) [5, 6], respectively, and a thermal noise amplitude of $(13.7\pm 0.3)\times 10^{-18}$ m Hz-1/2 (see Section 3.5).
Indeed, according to the fluctuation-dissipation theorem [13], the power spectral density of Brownian noise fluctuations at the coatings’ surface is proportional to the rate of elastic energy dissipation within the thin coatings’ layers. This rate, quantified by the loss angle $\varphi$, must hence be reduced as much as possible to decrease coating thermal noise. At room temperature, the loss angle of tantala and tantala-titania layers is about an order of magnitude higher than that of silica layers [10, 14], making it the dominant source of dissipation. So, since the early 2000s, a considerable research effort has been committed to developing alternative high-index coating materials of lower loss angle [15, 16, 17].
@@ 2.1 Samples
The thin films used in this work were deposited in a custom-developed facility at the LMA. We used several sets of deposition parameters, in order to test the hypothesis of a correlation between the synthesis conditions of the thin-film samples and their composition and optical and mechanical properties. Hence we varied the kinetic energy and current intensity of the sputtering particles, as well as the elemental composition of the ion beam impinging on the target. At the same time, for every deposition run, we kept the beam/target and target/substrate holder angles constant and equal to 45 deg [34, 35], the working pressure during deposition was of the order of $10^{-4}$ mbar, the substrates were heated up to a temperature between 100 and 150 C while rotating at a speed of 3 rpm. Prior to deposition, the base pressure inside the chamber was lower than $2\times 10^{-7}$ mbar.
@@
Table 1: Relevant deposition parameters of IBS SiNx thin-film samples: ion beam voltage $V$ and current $I$, Ar and N2 gas mass flows fed to the ion source, and average deposition rate $\dot{t}$. Uncertainties on $\dot{t}$ are $\leq 1$% (not shown).
@@
Figure 1: Average deposition rate $\dot{t}$ of IBS SiNx thin-film samples, as a function of ion beam settings: voltage $V$ and current $I$, and ratio N2/Ar of gas mass flows to the source. Uncertainties on $\dot{t}$ are $\leq 1$% (not shown).
@@ 2.2 Annealing
High-temperature thermal treatment, or annealing, is known to generally minimize the optical absorption and loss angle of thin-film coatings [10, 36], hence we studied the effect of annealing on the properties of our specific samples.
@@ 2.3.1 Spectroscopic photometry
Samples were deposited on NEGS1 fused-silica substrates (25.4 mm diameter, 6 mm thickness), then measured for transmittance in the 400–1400 nm wavelength range, at normal incidence, using a Cary 7000 spectroscopic photometer by Agilent.
@@ 2.3.3 Photo-thermal deflection (PTD)
Samples for absorption measurements were deposited on NEGS1 fused-silica substrates (25.4 mm diameter, 6 mm thickness).
@@ 2.4.1 Rutherford back-scattering spectrometry (RBS)
At SAFIR, we used samples grown on glassy carbon substrates ($25\times 25$ mm, 0.5 or 1 mm thickness). Relative atomic concentrations of the samples were deduced by integrating the peaks in the acquired spectra with the SIMNRA software [44, 45]. We also used SIMNRA to simulate a small background signal observed in the low-energy region of all spectra, and subtracted the results of those simulations from the oxygen and nitrogen contributions. We modeled that background signal as arising from 0.1 to 0.3% oxygen ambient contamination of the surface of the glassy carbon substrates.
@@ 2.5 Mechanical properties
We estimated the loss angle $\varphi$ of the samples by measuring the frequency and ring-down time of several mechanical resonances [46] of Corning 7980 fused-silica disks (50.0 mm diameter, 0.5 or 1.0 mm thickness), before and after coating deposition, in the 2.5-30 kHz range. In order to avoid systematic damping from suspension and ambient atmospheric pressure, we used a clamp-free, in-vacuum Gentle Nodal Suspension (GeNS) system [47]. More details about our experimental setup and data analysis can be found in our previous works [10, 48].
We derived the volumetric mass density $\rho$ from mass, area and thickness measurements. We measured the mass with an analytical balance, the area was estimated by measuring the diameter of our coated disks with a caliper, and the thickness was inferred from photometric and ellipsometric measurements.
@@ 2.6 Grazing-incidence X-ray diffraction (GI-XRD) and X-ray reflectivity (XRR)
For both GI-XRD and XRR measurements, samples were grown on single-side polished (100) silicon substrates (76.2 mm diameter, 0.5 mm thickness).
@@ 2.7 Ar desorption
As a consequence of the deposition process, the presence of argon in the samples was highly probable. In order to assess whether such contaminant was desorbed upon annealing, X-ray photoelectron spectroscopy and thermal-desorption spectroscopy measurements were performed on samples deposited on single-side polished (100) silicon wafers (76.2 mm diameter, 0.5 mm thickness).
@@ 2.7.2 Thermal-desorption spectroscopy (TDS)
TDS measurements were carried out by placing the samples in the quartz tube of an MTI GSL100 vacuum furnace. The tube was connected to a Pfeiffer Vacuum QMS200 mass spectrometer, its base pressure was of the order of 10-7 mbar. The furnace heating rate was set at 180 C/hour, and surveys were performed from ambient temperature up to the maximum temperature of 850 C allowed by the setup.
@@ 2.8 Coating Brownian thermal noise (CTN)
Samples for CTN measurements were deposited on Corning 7980 fused-silica substrates (25.4 mm diameter, 6 mm thickness).
@@ 3.1 Mitigation of water vapor contamination
In the early phases of this work, ever since the very first characterizations were performed, an evident anomaly was observed in the acquired photometric spectra. We interpreted it as a consequence of a gradient of refractive index $n$ along the direction of the samples’ thickness, with $n$ always decreasing from the surface towards the substrate. Under this hypothesis, a gradient as high as 27% was required to explain the measured data.
@@
Thus we performed a series of measurements of Fourier-transform infrared spectroscopy (FTIR), Raman spectroscopy, XPS (also coupled to sputtering along the direction of sample thickness, achieved via focused ion beam etching), RBS and ERDA. All those measurements, not shown here for conciseness, revealed an abundant presence of oxygen and hydrogen within the samples, whose relative atomic concentrations were as high as 34% and 5%, respectively. Incidentally, we were able to verify that the nitrogen-to-silicon ratio N/Si was equal to or less than 1.0 in all those samples, including those deposited using a pure N2 sputtering beam. This clearly indicated the presence of excess silicon, only partly oxidized, which may cause optical absorption.
@@ 3.2.1 Refractive index and extinction
We found that $n$ was a monotonic increasing function of the ion beam voltage $V$, in agreement with results found in the literature [18, 27], whereas we observed that the ion beam elemental composition and current $I$ only had a limited effect on it.
The voltage $V$ also had a noticeable effect on the $k$ coefficient, whose minimum was obtained at 700 V, and the same held for the ion beam composition. The lowest $k$ values were obtained for a mass flow ratio N2/Ar equal to 2. This outcome is in contrast with results reported in the literature, indicating that the optimal ratio would rather be 4 [54]. Regardless, the largest effect on $k$ was obtained by varying the ion beam current $I$. At $V=700$ V, we were able to decrease $k$ by more than an order of magnitude, from $2.1\times 10^{-4}$ to $1.5\times 10^{-5}$, by increasing $I$ up to 195 mA. That was the highest current value allowed by our experimental setup in order to have a stable plasma in the ion beam source.
Thus a minimum of extinction was achieved with sample 21025 for $V=700$ V, $I=195$ mA and N2/Ar = 2. Since then, we used those same parameters to deposit multiple nominally identical samples and observed repeatable results. For instance, the average extinction coefficient and refractive index of a subset of fifteen samples were $k=(1.5\pm 0.2)\times 10^{-5}$ and $n=2.04\pm 0.01$, respectively, where the uncertainties are standard deviations. After in-air annealing at 900 C for 10 hours, the extinction decreased to $k=(6.4\pm 0.2)\times 10^{-6}$. Compared to the tantala-titania layers in current mirror coatings of gravitational-wave interferometers, which show $n=2.09\pm 0.01$ and $k\sim 10^{-7}$ [10], the refractive index of sample 21025 is only 2% lower. Its extinction, however, while being to the best of our knowledge amongst the lowest ever measured on this class of thin films, is still excessively high. Indeed Wallace et al. measured significantly lower extinction values, although this was very likely due to the fact that their non-stoichiometric SiNx thin films were highly oxidized (oxynitrides) [22].
@@ 3.3 Elemental composition
In order to study more thoroughly the conditions that resulted in minimal optical absorption, we focused our attention on samples deposited on glassy carbon substrates, which allowed to clearly identify and discard the substrate background signal, and using $I=195$ mA, N2/Ar = 2 and varying $V$. Table 4 presents the relative atomic concentrations of samples 22020, 22022 and 22023, as deduced from acquired RBS spectra. Sample 22020 was deposited using $V=700$ V as in sample 21025, which featured the lowest optical absorption. Samples 22022 and 22023 were deposited using $V=500$ V and $V=900$ V, respectively. By way of example, Fig. 5 shows the measured and simulated RBS spectra of samples 22020 and 22023.
@@
Comparative analysis of RBS results of Table 4 first shows that using $I=195$ mA and N2/Ar = 2 settings slightly increased the N/Si ratio, which reached a value of about 1.1. The beam voltage $V$, on the other hand, only appreciably changed it for $V=900$ V, when it equalled 1.0. This seems to contradict the results found in the literature, which indicate that it would be possible to determine a range of $V$ values yielding stoichiometric samples [18, 25, 28, 29, 55]. It should be noted however that these results appear to be in sharp disagreement as to the values to be used. For instance, Bouchier et al. found that stoichiometric samples were obtained using $V\leq 500$ V [28, 29, 55], while Huang et al. and Lee et al. found that stoichiometric samples were instead obtained for 700 V $<V<$ 900 V [25] or even $V=1100$ V [18], respectively. Anyway, it remains that, in our specific case, the best deposition conditions identified so far produced sub-stoichiometric, nitrogen-poor samples.
By combining data of Tables 2 and 4 altogether, we observed that the refractive index $n$ at 1064 nm varies linearly with the Si/N ratios as measured by RBS. This is shown on Fig. 6. For the deposition conditions under discussion, we found that
@@ 3.3.1 Contaminants
Figure 7 shows the measured and simulated ERDA spectra of sample 23017, which are representative of those obtained from our other samples. It can be seen there that, at first glance, the measured relative concentration of hydrogen appeared to be highly inhomogeneous across the sample thickness. However, the two peaks observed at $\sim 453$ and $\sim 630$ keV were likely due to ambient contamination at the substrate and coating surface, respectively. Under this hypothesis, the hydrogen content of sample 23017 should be evaluated in the region in between the peaks, hence amounting to 0.4%. If the peaks were to be included in the analysis, the relative concentration would indeed increase, up to 1.2%. The small signal observed below 400 keV was identified as an artifact of the experimental setup and not considered in our analysis.
Figure 7: Measured and simulated ERDA spectra of IBS SiNx thin-film sample 23017, deposited on glassy carbon.
@@ 3.3.2 Ar desorption
Figure 8: Desorption of Ar from IBS SiNx thin-film sample 22027 upon in-vacuum annealing, compared to the background signal of the experimental setup. The heating rate was 180 C/h.